OpenStudy (anonymous):

A 45-45-90 triangle has a hypotenuse of length 7. What is the length of one of its legs? If necessary, round your answer to two decimal places.

OpenStudy (amistre64):

7 more than normal

OpenStudy (anonymous):

7/root2

OpenStudy (amistre64):

i asked me math teacher today why they call it a normal; he doesnt know :)

OpenStudy (anonymous):

Well: a^2+b^2=c^2. If the hypotenuse is 7 you have: a^2+b^2=49 BUT its a 45-45-90. So the legs are the same so: a^2+a^2=49 2a^2=49 a^2=49/2 a=7/sqrt(2)

OpenStudy (anonymous):

sqrt(49/2)=4.95

OpenStudy (anonymous):

lol amistre^^

OpenStudy (amistre64):

\[7*\frac{\sqrt{2}}{2}\]

OpenStudy (amistre64):

i gotta go to stats; have fun ;)

OpenStudy (anonymous):

Oh yeah?!?!? \[7e^{i \frac{\pi}{4}}-7i \sin(\frac{\pi}{4})!!\].

OpenStudy (amistre64):

now now... you keep up that notational language and your gonna poke out an "i"

OpenStudy (anonymous):

O.x